During the 2016 fiscal year, households in economy spent 75 per cent of their disposal income on consumption as well as 500 consumption expenditure which is independent of income. Total government expenditure which stood at 800 was supposed to be financed from a proportional tax levy of 20 per cent of national and VAT of 100. Total private investment spending was made up of 400 whereas export was 400 and an import of 500
a. Determine the equilibrium national income for this economy.
b. Determine the consumption and savings levels at equilibrium national income
1
Expert's answer
2018-08-03T15:46:08-0400
During the 2016 fiscal year, households in economy spent 75 per cent of their disposal income on consumption as well as 500 consumption expenditure which is independent of income. Total government expenditure which stood at 800 was supposed to be financed from a proportional tax levy of 20 per cent of national and VAT of 100. Total private investment spending was made up of 400 whereas export was 400 and an import of 500
a. The equilibrium national income for this economy is:
C = C0 + c*(Y - T),
NX = E - I,
Y = C + I + G + NX = 500 + 0.75*(Y - 0.2*Y - 100) + 400 + 800 + (400 - 500),
Y = 1525 + 0.6Y,
0.4Y = 1525,
Y = 3812.5.
b. The consumption is: C = C0 + c*(Y - T) = 500 + 0.75*(0.8*3812.5 - 100) = 2712.5.
Savings level at equilibrium national income is: S = (1 - c)*Y = 0.25*3812.5 = 953.125.